OpenStudy (anonymous):

evaluate 3x-2y for x= 1/4 and y = -1/3

OpenStudy (anonymous):

just substitute the values of x=1/4 and y=-1/3 into the equation.

OpenStudy (anonymous):

3(1/4) - 2(-1/3) and what do you get?

OpenStudy (anonymous):

17/12

OpenStudy (anonymous):

yes, that's correct.

OpenStudy (anonymous):

thank u so much i have done this

OpenStudy (anonymous):

so many times